### v1

# Dirac Sources for Nonmetricity and Torsion in Metric-affine Gravity

Abstract
Metric-affine gravity (GL(4) gauge theory) in 4-dimensions is coupled to a spacetime Dirac source field using the isomorphisms of the Lie algebra gl(4) to the Clifford algebras Cl(3,1) and Cl(2,2). A simple transformation relates the generators of Cl(3,1) to a real representation of Cl(2,2), while the real representation of Cl(2,2) serves directly as a basis for the Lie algebra gl(4). Therefore, although GL(4) does not contain a spinor representation of the Lorentz group, expanding its Lie algebra in the Cl(2,2) basis gives a Clifford valued connection with well-defined coupling to Dirac spinors. Variation of the expansion coefficients gives new Dirac sources for both torsion and nonmetricity, separated by identifying the so(3,1) basis within the gl(4) basis.
Keywords: General linear gauge theory, GL(4) gravity, Poincaré gauge theory, torsion, nonmetricity, non-metricity, nonmetric gravity, non-metric gravity, general relativity, metric-affine gravity, hypermomentum, particle-antiparticle asymmetry
James T Wheeler, Utah State University Department of Physics, 4415 Old Main Hill, Logan, UT 84322-4415, jim.wheeler@usu.edu
1 Introduction
When the metric and connection on a 4-dimensional manifold $\left(\mathcal{M},g,\Sigma\right)$ are treated independently, the connection may be divided by symmetry into nonmetricity and torsion, with the symmetric nonmetric part first explored by H. Weyl [1, 2, 3, 4] and the antisymmetric torsion part introduced by É. Cartan [5, 6, 7, 8, 9]. These tensors may both be defined as covariant exterior derivatives
$\mathbf{Q}_{ab}$ $=$ $\boldsymbol{\mathfrak{D}}g_{ab}\>=\>\mathbf{d}g_{ab}-g_{cb}\boldsymbol{\Sigma}_{\;\;\;a}^{c}-g_{ac}\boldsymbol{\Sigma}_{\;\;\;b}^{c}$
$\boldsymbol{\mathfrak{T}}^{a}$ $=$ $\boldsymbol{\mathfrak{D}}\mathbf{e}^{a}\>=\>\mathbf{d}\mathbf{e}^{a}-\mathbf{e}^{b}\land\boldsymbol{\Sigma}_{\;\;\;b}^{a}$
where the basis 1-form $\mathbf{e}^{a}$ and the metric $g_{ab}$ are related by the inner product $\left\langle\mathbf{e}^{a},\mathbf{e}^{b}\right\rangle=g^{ab}$ with $g^{ab}g_{bc}=\delta_{c}^{a}$. Here $\boldsymbol{\mathfrak{D}}$ is the covariant exterior derivative with connection 1-form $\boldsymbol{\Sigma}_{\;\;\;b}^{a}$, the 2-form $\boldsymbol{\mathfrak{T}}^{a}$ is the torsion, and the 1-form $\mathbf{Q}_{ab}$ is the nonmetricity. If the basis $\mathbf{e}^{a}$ is orthonormal $\left\langle\mathbf{e}^{a},\mathbf{e}^{b}\right\rangle=\eta^{ab}$ the nonmetricity becomes algebraic
$\mathbf{Q}_{ab}$ $=$ $\cancel{\mathbf{d}\eta_{ab}}-\eta_{cb}\boldsymbol{\Sigma}_{\;\;\;a}^{c}-\eta_{ac}\boldsymbol{\Sigma}_{\;\;\;b}^{c}=\>-\left(\boldsymbol{\Sigma}_{ba}+\boldsymbol{\Sigma}_{ab}\right)$
Alternatively, if we use a coordinate basis $\mathbf{e}^{a}=\delta_{\mu}^{\;\;\;a}\mathbf{d}x^{\mu}$ then the torsion becomes algebraic
$\boldsymbol{\mathfrak{T}}^{a}$ $=$ $\cancel{\mathbf{d}^{2}x^{a}}-\mathbf{e}^{b}\land\boldsymbol{\Sigma}_{\;\;\;b}^{a}\>=-\Sigma_{\;\;\;\mu\nu}^{a}\mathbf{d}x^{\mu}\land\mathbf{d}x^{\nu}$
so that $\mathfrak{T}_{\;\;\;\mu\nu}^{a}=\Sigma_{\;\;\;\nu\mu}^{a}-\Sigma_{\;\;\;\mu\nu}^{a}$. Here we use Latin indices for orthonormal frames and Greek indices for coordinate frames.
Torsion arises naturally within ECSK gravity, including Cartan’s introductory work [5, 6, 7, 8, 9] together with [10, 11, 12, 13, 14, 15, 16, 17, 18]. This may be understood as the gauge theory of the Poincaré group using the Einstein-Hilbert form of the action but with an asymmetric connection. In ECSK gravity the curvature is the field strength of the Lorentz connection and the torsion is the field strength of the gauge field of translations.
Nonmetric geometries include Weyl geometry (see review in [19]), which depends on a single vector given by the trace of the nonmetricity. Geometries with general nonmetricity have been largely confined to metric-affine gravity [20, 21, 22, 23, 24, 25], the gauge theory of the general linear group $GL\left(4,\mathbb{R}\right)$  [^1] But see [26], showing a relationship between nonmetricity and the conformal curvatures.. The metric and connection are necessarily treated independently since the general linear group does not single out a metric.
Although neither torsion nor nonmetricity has any direct experimental support, the study of these more general geometries gives a broader arena for tests of general relativity, and may lead to their ultimate measurement or explain why we do not see them. Of particular interest in this regard are possible Standard Model sources. As we show in detail in the next Section, neither torsion nor nonmetricity is driven by Yang-Mills gauge theories or scalars, because the actions of those fields do not depend on the spacetime connection. Dirac and Rarita-Schwinger spinor fields, however, do provide sources for torsion in ECSK gravity [27].
The goal for our current work is to identify Dirac sources for both torsion and nonmetricity in $GL\left(4\right)$ gravity. It is important to observe that since the sources for torsion and nonmetricity both depend on the overall symmetry of the connection, the source for torsion differs between ECSK and $GL\left(4\right)$ (metric-affine) gravity. Concretely, Poincaré gauge theory couples only the totally antisymmetric part of the torsion to a single Dirac pseudo-current, but as we show here, all 16 Dirac currents play a role in metric-affine gravity. Since $GL\left(4\right)$ does not have a preferred metric, the sources for torsion differ strongly from the predictions of ECSK theory. The full range of Dirac sources for generic nonmetricity have not previously been identified.
The principal issue encountered with Dirac sources in metric-affine gravity is that the general linear group $GL\left(4,\mathbb{R}\right)$ does not have finite dimensional spinor representations. Early studies of metric-affine gravity [20] noted that the source for the metric-affine connection, called hypermomentum, does not easily include spinors:
“This last remark [that the traceless nonmetricity depends on generators of general linear transformations] seems to preclude the definition of a canonical hypermomentum tensor for spinor fields, since $GL\left(4,\mathbb{R}\right)$ has no spinor representations. For the time being, at least, we exclude spinor fields from our consideration . . ..”
Since the Yang-Mills and scalar fields of the standard model do not provide sources for nonmetricity, excluding spinor sources ignores direct coupling of nonmetricity to the Standard Model. Nonetheless, coupling is possible because the connection depends only on the Lie algebra $\mathfrak{gl}\left(4\right)$, which does have spinor representations. An interesting recent study [25] includes Dirac sources, but unlike our presentation restricts the fermion coupling to the antisymmetric part of the connection. This restriction means that [25] only couples spinors to traces of the torsion and nonmetricity. Our current work couples the full connection to spinors.
[…]
A second seeming obstacle to finding spinor sources in $GL\left(4\right)$ gravity is the definition of the spinors themselves, because spinors are defined as representations of a Clifford algebra which in turn depends on the Lorentz metric $\eta_{ab}$ through the definition of the gamma matrices, $\left\{\gamma^{a},\gamma^{b}\right\}=-2\eta^{ab}1$. The resolution lies in the independence of metric and connection in metric-affine gravity, which allows us to choose the Minkowski metric. Since any metric we choose will be not be compatible with the connection, we have this freedom.
A class of sources for the nonmetricity in metric-affine gravity (the gauge theory of the general linear group), is defined by the variation of the action with respect to the connection [18]
$e\Delta_{i}^{\;\;\;jk}\equiv-\frac{\delta\mathcal{L}_{source}}{\delta\Sigma_{\;\;\;jk}^{i}}$
where the metric and connection are independent. The source tensor $\Delta_{i}^{\;\;\;jk}$ is called the hypermomentum. By expanding the $GL\left(4\right)$ connection $\Sigma_{\;\;\;jk}^{i}$ in a $\mathfrak{cl}\left(3,1\right)$ basis we can describe the hypermomentum of the spinor field. The process is simplified by working with the Clifford algebra $\mathfrak{Cl}\left(2,2\right)$ built from the basis gamma matrices of $\mathfrak{spin}\left(2,2\right)$, which has real representations. Since $\mathfrak{Cl}\left(2,2\right)$ has real representations, a real basis for $\mathfrak{Cl}\left(2,2\right)$ may serve directly as a basis for $\mathfrak{gl}\left(4\right)$. At the same time there is a simple transformation from that real basis to a basis for $\mathfrak{Cl}\left(3,1\right)$.
We use the Einstein-Hilbert form of the gravitational action, but build the curvature from the general linear connection, leading to both torsion and nonmetric dependence. Then, to separate the torsion and nonmetric dependence of the action, we identify the $SO\left(3,1\right)$ subgroup of $GL\left(4\right)$. The remaining dependence is identified with nonmetricity. For sources we consider the three types of matter field within the Standard Model: the Higgs scalar doublet, $U\left(1\right)$ and Yang-Mills fields, and Dirac spinors. We show in the next Section that only Dirac fields provide sources  [^2] It is easy to see that non-gauge tensor fields can provide sources. Let $S_{T}=\int\left(D^{\alpha}V^{\mu\ldots\nu}D_{\alpha}V_{\mu\ldots\nu}+m^{2}V^{\mu\ldots\nu}V_{\mu\ldots\nu}\right)\sqrt{-g}d^{4}x$ and vary $H_{\beta\mu\nu}=-\frac{1}{2}\left(Q_{\beta\mu\nu}+Q_{\beta\nu\mu}-Q_{\mu\nu\beta}\right)$. With $\delta_{H}\left(D_{\alpha}V_{\mu\ldots\nu}\right)=-V_{\beta\ldots\nu}\delta H_{\;\;\;\mu\alpha}^{\beta}-\cdots-V_{\mu\ldots\beta}\delta H_{\;\;\;\nu\alpha}^{\beta}$we have a coupling $\delta_{H}S_{T}=-2\int D^{\alpha}V^{\mu\ldots\nu}\left(V_{\beta\ldots\nu}\delta_{\mu}^{\rho}+\cdots+V_{\mu\ldots\beta}\delta_{\nu}^{\rho}\right)\delta H_{\;\;\;\rho\alpha}^{\beta}\sqrt{-g}d^{4}x$ so that the nonmetricity has a source. This dependence vanishes with antisymmetrization so that Yang-Mills fields depend only on the internal $SU\left(N\right)$ connection..
In the next Section we expand the metric-affine curvature, separating the torsion and nonmetric dependence from the usual Riemannian curvature, then vary to find the gravitational contribution to the torsion and nonmetric field equations. We show the well-known result that in vacuum the torsion vanishes; for the nonmetric variation we find that only the totally symmetric part of the nonmetricity vanishes. The remaining part of the nonmetricity consists of the Weyl vector and an antisymmetric piece which may be absorbed into the torsion [26]. In Section 3 we present details of the Clifford algebras $\mathfrak{Cl}\left(3,1\right)$ and $\mathfrak{Cl}\left(2,2\right)$, then write the coupled form of the Dirac action. In Section 4 we separate the generators of the $SO\left(3,1\right)$ subgroup of $GL\left(4\right)$, allowing us to identify the part of the connection variation relating to torsion. Our main results, the Dirac sources for torsion and nonmetricity in metric-affine gravity, are presented in Section 5, with some simple cases in the short final Section.
For the orthonormal vector basis we use lower case Latin letters, $\eta_{ab},\mathbf{e}^{a}$. For coordinates, if needed, we use lower case Greek, $g_{\alpha\beta},v^{\mu}$. For spinors we use upper case Latin indices, e.g., $\psi^{A},\left[\gamma^{a}\right]_{\;\;\;C}^{B}$, and for the Clifford basis we use upper case Greek, $\Gamma^{\Delta}\in\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\}$. The Clifford basis may be divided into symmetric $\Gamma^{\Delta_{s}}$ and antisymmetric $\Gamma^{\Delta_{a}}$ parts. When needed we use the Dirac form of the gamma matrices, given in 3.
2 Einstein-Hilbert action with torsion and nonmetric additions
2.1 Decomposition of the connection and curvature
The spacetime arena we consider is the class of geometries modeled on the principle $GL\left(4\right)$ fiber bundle based on the quotient of the affine group $A\left(4\right)=R^{4}\rtimes GL\left(4\right)$ by the general linear group, $\mathcal{M}_{0}^{4}=A\left(4\right)/GL\left(4\right)$. Generalizing the homogeneous quotient manifold $\mathcal{M}_{0}^{4}\rightarrow\mathcal{M}^{4}$ and the flat Maurer-Cartan connection to $\boldsymbol{\Sigma}_{\;\;\;b}^{a}$, we have the Cartan structure equations:
$\mathbf{d}\mathbf{e}^{a}$ $=$ $\mathbf{e}^{b}\land\boldsymbol{\Sigma}_{\;\;\;b}^{a}+\boldsymbol{\mathfrak{T}}^{a}$ (1)
[…]
$\mathbf{Q}_{ab}\>=\;\boldsymbol{\mathfrak{D}}g_{ab}$ $=$ $\mathbf{d}g_{ab}-g_{cb}\boldsymbol{\Sigma}_{\;\;\;a}^{c}-g_{ac}\boldsymbol{\Sigma}_{\;\;\;b}^{c}$ (3)
[…]
To include Standard Model sources we extend this spacetime background by including the Standard Model symmetries. The quotient $\mathcal{M}_{0}^{4}=\left[A\left(4\right)\times U\left(1\right)\times SU\left(2\right)\times SU\left(3\right)\right]/\left[GL\left(4\right)\times U\left(1\right)\times SU\left(2\right)\times SU\left(3\right)\right]$ is then a principal fiber bundle with base manifold $\mathcal{M}_{0}^{4}$ and fibers $GL\left(4\right)\times U\left(1\right)\times SU\left(2\right)\times SU\left(3\right)$. In addition to Eqs.(1) and (2), the Cartan equations now include
$\mathbf{d}\mathbf{A}$ $=$ $\mathbf{F}$
$\mathbf{d}\mathbf{W}^{a}$ $=$ $-\frac{1}{2}c_{\;\;\;bc}^{a}\mathbf{W}^{b}\wedge\mathbf{W}^{c}+\mathbf{F}^{a}\quad\left(a,b,c=1,2,3\right)$
$\mathbf{d}\mathbf{B}^{K}$ $=$ $-\frac{1}{2}f_{\;\;\;LM}^{K}\mathbf{B}^{L}\wedge\mathbf{B}^{M}+\mathbf{H}^{K}\quad\left(K,L,M=1,\ldots,8\right)$ (4)
[…]
Choose the orthonormal basis, $g_{ab}=\eta_{ab}$, so that $\boldsymbol{\Sigma}_{ba}+\boldsymbol{\Sigma}_{ab}=-\mathbf{Q}_{ab}$ and let $\boldsymbol{\Omega}_{ab}=\boldsymbol{\Sigma}_{\left[ab\right]}$. Then the connection is
$\boldsymbol{\Sigma}_{ab}$ $=$ $-\frac{1}{2}\mathbf{Q}_{ab}+\boldsymbol{\Omega}_{ab}$ (5)
Substituting Eq.(5) into Eq.(1) results in
$\mathbf{d}\mathbf{e}^{a}$ $=$ $\mathbf{e}^{b}\land\boldsymbol{\Omega}_{\;\;\;b}^{a}+\boldsymbol{\mathfrak{T}}^{a}-\mathbf{Q}^{a}$
where we define the nonmetric 2-form $\mathbf{Q}^{a}\equiv\frac{1}{2}\mathbf{e}^{b}\land\mathbf{Q}_{\;\;\;b}^{a}$. We may eliminate explicit dependence on the nonmetric 2-form $\mathbf{Q}^{a}$ by the field redefinition, $\mathbf{T}^{a}\equiv\boldsymbol{\mathfrak{T}}^{a}-\mathbf{Q}^{a}$, leaving the nonmetricity determined by its remaining, totally symmetric part $Q_{\left(abc\right)}$. The resulting structure equation determines a Lorentz connection with torsion
$\mathbf{d}\mathbf{e}^{a}$ $=$ $\mathbf{e}^{b}\land\boldsymbol{\Omega}_{\;\;\;b}^{a}+\mathbf{T}^{a}$
We solve for $\boldsymbol{\Omega}_{\;\;\;b}^{a}$ by defining the torsion-free, metric-compatible spin connection $\boldsymbol{\omega}_{\;\;\;b}^{a}$ satisfying $\mathbf{d}\mathbf{e}^{a}=\mathbf{e}^{b}\land\boldsymbol{\omega}_{\;\;\;b}^{a}$. Then setting $\boldsymbol{\Omega}_{\;\;\;b}^{a}=\boldsymbol{\omega}_{\;\;\;b}^{a}-\mathbf{C}_{\;\;\;b}^{a}$, the contorsion $\mathbf{C}_{\;\;\;b}^{a}$ must satisfy $0=\mathbf{e}^{b}\land\mathbf{e}^{c}\left(-C_{\;\;\;bc}^{a}+\frac{1}{2}T_{\;\;\;bc}^{a}\right)$. Lowering $a$ and cycling indices, we add the first two permutations and subtract the third to find
$C_{abc}$ $=$ $\frac{1}{2}\left(T_{abc}+T_{bca}-T_{cab}\right)$ (6)
We may recover the torsion as $\mathbf{e}^{b}\wedge\mathbf{C}_{\;\;\;b}^{a}=\mathbf{T}^{a}$.
With $\boldsymbol{\Omega}_{\;\;\;b}^{a}=\boldsymbol{\omega}_{\;\;\;b}^{a}-\mathbf{C}_{\;\;\;b}^{a}$, Eq.(5) becomes $\boldsymbol{\Sigma}_{\;\;\;b}^{a}=\boldsymbol{\omega}_{\;\;\;b}^{a}-\mathbf{C}_{\;\;\;b}^{a}-\frac{1}{2}\mathbf{Q}_{\;\;\;b}^{a}$. Substituting this into Eq.(2) results in
$\boldsymbol{\mathfrak{R}}^{ab} =\displaystyle= 𝓡$$\boldsymbol{\mathcal{R}}^{ab}-\frac{1}{2}\mathbf{D}\mathbf{Q}^{ab}-\frac{1}{4}\mathbf{Q}^{cb}\land\mathbf{Q}_{\;\;\;c}^{a}-\mathbf{Q}^{c(a}\land\mathbf{C}_{\;\;\;c}^{b)}$ (7)
[…]
2.2 The action
[…]
$\delta_{\Omega}S_{Grav}$ $=$ $\kappa\int\delta\boldsymbol{\Omega}^{eb}\land\mathbf{C}_{\;\;\;e}^{a}\wedge\mathbf{e}^{c}\wedge\mathbf{e}^{d}e_{abcd}$
$=$ $\kappa\int\delta\Omega_{\quad f}^{eb}\left(C_{\;\;\;eb}^{f}-C_{\;\;\;be}^{f}-C_{\;\;\;ea}^{a}\delta_{b}^{f}+C_{\;\;\;ba}^{a}\delta_{e}^{f}\right)\boldsymbol{\Phi}$
where the volume form $\boldsymbol{\Phi}$ is given by the Hodge dual of one, $^{*}1$. Using Eq.(6) the integrand $\mathscr{T}_{\;\;\;eb}^{g}=C_{\;\;\;eb}^{g}-C_{\;\;\;be}^{g}-C_{\;\;\;ea}^{a}\delta_{b}^{g}+C_{\;\;\;ba}^{a}\delta_{e}^{g}$ reduces to a trace-altered form of the torsion
$\mathscr{T}_{\;\;\;bc}^{a}=T_{\;\;\;bc}^{a}-\delta_{b}^{a}T_{\;\;\;ec}^{e}+\delta_{c}^{a}T_{\;\;\;eb}^{e}$ (9)
Equate this to the source contribution,
$\kappa\mathscr{T}_{\;\;\;bc}^{a}$ $=$ $-\frac{\delta\mathcal{L}_{Source}}{\delta\Omega^{abc}}$ (10)
The variation of nonmetricity is similar. We find
$\delta_{Q}S_{Grav}$ $=$ $-\frac{\kappa}{4}\int\delta_{Q}Q_{\quad f}^{eb}\left(Q_{\;\;\;eb}^{f}+Q_{\;\;\;be}^{f}+Q_{\;\;\;ea}^{a}\delta_{b}^{f}+Q_{\;\;\;ba}^{a}\delta_{e}^{f}\right)\boldsymbol{\Phi}$
Defining the trace-altered nonmetricity,
$\mathcal{\mathscr{Q}}_{acb}$ $\equiv$ $-\frac{1}{4}\left(Q_{acb}+Q_{abc}-Q_{\;\;\;be}^{e}\eta_{ca}-Q_{\;\;\;ce}^{e}\eta_{ba}\right)$ (11)
and including sources, the field equation is
$\mathcal{\mathscr{Q}}_{cba}$ $=$ $\frac{4}{\kappa}\frac{\delta\mathcal{L}_{Source}}{\delta Q^{abc}}$ (12)
The source term on the right has been called the hypermomentum [17, 24, 28]. Notice that we may write $\mathcal{\mathscr{Q}}_{acb}$ as $\mathcal{\mathscr{Q}}_{acb}=-\left(\frac{1}{2}Q_{a\left(bc\right)}-\frac{1}{4}\eta^{de}\left(Q_{b\left(de\right)}\eta_{ca}+Q_{c\left(de\right)}\eta_{ba}\right)\right)$ so that the nonmetric field equation is independent of the nonmetric 2-form $\mathbf{Q}_{a}$ which has been absorbed into the torsion.
We find it helpful to write the variation in the general form
$\delta S_{Grav}$ $=$ $\kappa\int\left(\delta\Omega_{\quad c}^{ab}\mathscr{T}_{\;\;\;ab}^{c}-\frac{1}{4}\delta Q_{\quad a}^{bc}\mathscr{Q}_{\;\;\;bc}^{a}\right)\boldsymbol{\Phi}$ (13)
where $\Omega_{\quad c}^{ab}$ and $Q_{\quad a}^{bc}$ are the antisymmetric and symmetric parts of the connection, respectively.
In vacuum, $\mathscr{T}_{\;\;\;bc}^{a}=0$ and $\mathcal{\mathscr{Q}}_{cba}=0$. The torsion therefore vanishes, while contraction of $\mathcal{\mathscr{Q}}_{cba}=0$ shows that $Q_{\;\;\;ae}^{e}=\frac{1}{4}Q_{\;\;\;ea}^{e}=2W_{a}$ where $W_{b}$ is the Weyl vector. It follows that $Q_{c\left(ab\right)}=2\eta_{c(a}W_{b)}$ so that the full solution for the nonmetricity in vacuum is $Q_{abc}=2\eta_{c(a}W_{b)}+Q_{a\left[bc\right]}$ and we may write $\mathbf{Q}_{ab}$ in terms of $\boldsymbol{\omega}$ and $\mathbf{Q}^{a}$ only, both of which may be absorbed into a Weyl geometry with torsion [26].
3 Clifford basis and the metric-affine interaction
The vacuum Dirac action is
$S_{D,V}$ $=$ $\alpha\int\psi^{\dagger}h\left(i\cancel{\partial}-m\right)\psi\>d^{4}x$ (14)
where $\psi\in\mathbb{C}^{4}$ is a Dirac spinor, $h$ is the Hermitian metric $\left\langle\chi,\psi\right\rangle=\chi^{\dagger}h\psi\equiv\bar{\chi}\psi$, and $\cancel{\partial}=\gamma^{a}\partial_{a}$ where the $\gamma^{a}$ are four $4\times 4$ matrices satisfying
$\left\{\gamma^{a},\gamma^{b}\right\}=-2\eta^{ab}\mathbf{1}$ (15)
with $a,b,\ldots=0,1,2,3$. The commutators $\sigma^{ab}=\frac{1}{2}\left[\gamma^{a},\gamma^{b}\right]$ span the Lie algebra of the Lorentz group, so that $\Lambda\left(w^{ab}\right)=\exp\left(\frac{1}{2}w_{ab}\sigma^{ab}\right)$ are Lorentz transformations. The action is real, $S_{D}^{*}=S_{D}$, and the Hermitian metric satisfies $\gamma^{a\dagger}h=h\gamma^{a}$.
In nonflat spacetimes, the Dirac equation requires a covariant derivative. The connection becomes a Lie-algebra valued 1-form $\boldsymbol{\beta}_{A}G^{A}$ where generators $G^{A}$ form a basis for a spinor representation of the Lie algebra of the gravitational symmetry. In spaces with compatible $SO\left(p,q\right)$ invariant metric $\eta_{ab}$, this takes the form $\frac{1}{2}\boldsymbol{\omega}_{ab}\sigma^{ab}$ where the coefficients are characterized by antisymmetry $\boldsymbol{\omega}_{ab}=-\boldsymbol{\omega}_{ba}$. The $\left(p,q\right)$ signature then enters when an index is raised, $\boldsymbol{\omega}_{\;\;\;b}^{a}=\eta^{ac}\boldsymbol{\omega}_{cb}$. In particular this applies to a Lorentz connection, precluding any nonmetricity $\mathbf{Q}_{ab}$ since in an orthonormal basis nonmetricity depends entirely on the symmetric part $\mathbf{Q}_{ab}=-\boldsymbol{\omega}_{ab}-\boldsymbol{\omega}_{ba}$. Therefore, nonmetricity is exactly the Lorentz- or $SO\left(p,q\right)$-violating part of the spacetime connection.
The requirement for a spinor representation requires generators $G^{A}$ that act linearly on spinors. A complete set of such operators is given by the Clifford algebra associated with a spin group. We can accomplish this for a general linear connection by writing a $\mathfrak{gl}\left(4\right)$ basis as elements of either $\mathfrak{Cl}\left(3,1\right)$ or $\mathfrak{Cl}\left(2,2\right)$.
We describe these Clifford algebras here, then identify the resulting spinor-$\mathfrak{gl}\left(4\right)$ interaction.
3.1 Cl(3,1)
The Clifford algebra $\mathfrak{Cl}\left(3,1\right)$ is the quotient of the free algebra of the $\gamma$-matrices by the symmetric relation (15). Equation (15) allows us to reduce any further symmetric products, so that a basis for the Clifford algebra is given by all antisymmetric products $\gamma^{a},\gamma^{[a}\gamma^{b]},\gamma^{[a}\gamma^{b}\gamma^{c]},\gamma^{[a}\gamma^{b}\gamma^{c}\gamma^{d]}$. Appending the identity, these antisymmetric products form a complete basis. They are generally written in the more convenient form
$\Gamma^{\Delta}$ $=$ $\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{d},\gamma_{5}\right\}$ (16)
where $\sigma^{ab}=\frac{1}{2}\left[\gamma^{a},\gamma^{b}\right],\>\gamma_{5}\gamma^{d}=\frac{i}{3!}e_{\;\;\;abc}^{d}\gamma^{a}\gamma^{b}\gamma^{c}$ and $\gamma_{5}=\frac{i}{4!}\varepsilon_{abcd}\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}$. Upper case Greek indices run $\Delta,\Omega,\ldots=1,\ldots,16$. The 16 matrices $\Gamma^{\Delta}$ span all complex $4\times 4$ matrices and satisfy the orthonormality relation
$\frac{1}{4}tr\left(\Gamma^{\Delta}\Gamma^{\Sigma\dagger}\right)=\delta^{\Delta\Sigma}$ (17)
Their completeness is central to our discussion, because it means we can expand the real, $4\times 4$ representation of the general linear Lie algebra as linear combinations $\beta_{\Delta}\Gamma^{\Delta}$. These linear combinations combine with spinor fields to form source currents, $\bar{\psi}_{A}\left[\beta_{\Delta}\Gamma^{\Delta}\right]_{\;\;\;B}^{A}\psi^{B}$.
To make the couplings of Dirac fields to nonmetricity explicit, we use the Dirac representation, $\gamma^{a}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, and $\gamma^{i}=\left(\begin{array}[]{cc}0&\sigma^{i}\\ -\sigma^{i}&0\end{array}\right)$. In this representation $\left[h\right]_{AB}=\left[\gamma^{0}\right]_{\;\;\;B}^{A}$, so the Hermitian inner product $\left\langle\chi,\psi\right\rangle=\chi^{\dagger}h\psi\equiv\bar{\chi}\psi$ we have $\bar{\chi}=\chi^{\dagger}h=\chi^{\dagger}\gamma^{0}$ in the usual way. In this representation, the remainder of the Clifford basis $\Gamma^{A}=\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\}$ is
$\begin{array}[]{ccccccccccccc}\sigma^{ij}&=&-i\varepsilon_{\;\quad k}^{ij}\left(\begin{array}[]{cc}\sigma^{k}\\ &\sigma^{k}\end{array}\right)&&&\sigma^{0i}&=&\left(\begin{array}[]{cc}&\sigma^{i}\\ \sigma^{i}\end{array}\right)&&&\gamma_{5}\gamma^{0}&=&\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)\\ \gamma_{5}\gamma^{i}&=&\left(\begin{array}[]{cc}-\sigma^{i}&0\\ 0&\sigma^{i}\end{array}\right)&&&\gamma_{5}&=&\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\end{array}$
Here $\sigma^{i}$ are the Pauli matrices and $1$ is the $2\times 2$ identity. Manipulation of the gamma matrices is familiar.
While complex linear combinations $\beta_{\Delta}\Gamma^{\Delta}$ can provide the real basis required for $\mathfrak{gl}\left(4\right)$, it is simpler to make use of the Clifford algebra associated with $\mathfrak{spin}\left(2,2\right)$ since this admits a real basis directly. The Clifford algebra $Cl\left(2,2\right)$ is simply related to $Cl\left(3,1\right)$ making it straightforward to characterize the contribution of Dirac spinors.
3.2 Cl(2,2)
Unlike $\mathfrak{spin}\left(3,1\right)$, the Lie algebra $\mathfrak{spin}\left(2,2\right)$ admits a real representation. This has the advantage of directly providing a basis for $\mathfrak{gl}\left(4,\mathbb{R}\right)$, while making only a slight change to the $\mathfrak{spin}\left(3,1\right)$ basis $\Gamma^{\Delta}$. As above, lower case Latin indices $a,b,\ldots=0,1,2,3,$ refer to either $SO\left(3,1\right)$ or $SO\left(2,2\right)$, while upper case Latin are spinor indices. It is convenient (but unusual!) to let $i,j,\ldots=0,1,3$.
A real form of gamma matrices for $\mathfrak{spin}\left(2,2\right)$ be chosen by inserting a single factor of $i$ on $\gamma^{2}$:
$\hat{\gamma}^{a}$ $\equiv$ $\left(\gamma^{0},\gamma^{1},i\gamma^{2},\gamma^{3}\right)$
These satisfy $\left\{\hat{\gamma}^{a},\hat{\gamma}^{b}\right\}=-2\hat{\eta}^{ab}$ where the metric is $\hat{\eta}_{ab}=diag\left(-1,1,-1,1\right)$. We confirm that $\hat{\sigma}^{AB},\hat{\gamma}_{5}\hat{\gamma}^{A}$ and $\hat{\gamma}_{5}=\gamma_{5}$ are all real, with $\hat{\gamma}_{5}\hat{\gamma}^{A}=\gamma_{5}\hat{\gamma}^{A}$ and
$\hat{\sigma}^{AB}$ $=$ $\frac{1}{2}\left[\hat{\gamma}^{A},\hat{\gamma}^{B}\right]=\left(\sigma^{01},i\sigma^{02},\sigma^{03},i\sigma^{12},i\sigma^{23},\sigma^{31}\right)$ (21)
Because the matrices $\hat{\Gamma}^{\Delta}=\left\{1,\hat{\gamma}^{A},\hat{\sigma}^{AB},\hat{\gamma}_{5}\hat{\gamma}^{A},\hat{\gamma}_{5}\right\}$ are all real, we may write the $\mathfrak{gl}\left(4,\mathbb{R}\right)$ connection as $\boldsymbol{\Sigma}=\mathbf{b}_{\Delta}\hat{\Gamma}^{\Delta}$ with the 1-forms $\mathbf{b}_{\Delta}$ real, $\Delta=1,\ldots,16$. Because we express $\hat{\Gamma}^{\Delta}$ in terms of $\Gamma^{\Delta}$ we easily find the action on Dirac spinors.
3.3 The covariant Dirac equation
Our action is $S=S_{Grav}+S_{D}$ where $S_{Grav}$ is given by Eq.(8). The Dirac action $S_{D}$ is adapted to metric-affine geometry by replacing the gradient with a covariant derivative, $\cancel{\partial}=\gamma^{a}e_{a}^{\;\;\;\mu}\partial_{\mu}\rightarrow\cancel{D}=\gamma^{a}e_{a}^{\;\;\;\mu}D_{\mu}$ where
$D_{\mu}\psi=\partial_{\mu}\psi-b_{A\mu}\hat{\Gamma}^{A}\psi$
With $b_{A\mu}$ real, $b_{A\mu}\hat{\Gamma}^{A}$ is the $GL\left(4\right)$ connection.
The spinor action must be made manifestly real, so we separate the real vacuum terms 14 from the interaction, $S_{D}=S_{D,V}-\alpha Re\int b_{aA}\psi^{\dagger}ih\gamma^{a}\hat{\Gamma}^{A}\psi$ to identify the interaction
$S_{interaction}$ $=$ $-\alpha Re\int b_{aA}\psi^{\dagger}ih\gamma^{a}\hat{\Gamma}^{A}\psi$ (22)
The contributions to the field equation are found by varying the real coefficients $b_{aA}$, but in this form it is not clear which terms will be sources for torsion and which will drive nonmetricity. To determine this, in the next Section we find the generators of the $SO\left(3,1\right)$ subgroup of $GL\left(4\right)$ in terms of the $\left[\hat{\Gamma}\right]_{\;\;\;B}^{A}$.
4 Nonmetricity vs. torsion: separating the SO(3,1) subgroup in the Clifford basis
To apply Eqs.(10) and (12) with the Dirac sources of Eq.(22) we expand the connection in the $\mathfrak{spin}\left(2,2\right)$ basis $\mathbf{b}_{A}\hat{\Gamma}^{A}$. To correctly interpret the results, we need to distinguish the torsion and nonmetric parts of the connection. Among the $\hat{\Gamma}^{\Delta}$ there must be real combinations that generate the real vector representation of the Lorentz group $SO\left(3,1\right)$. Varying the $\mathfrak{so}\left(3,1\right)$ generators will give the coupling to torsion, with the remaining independent combinations giving the field equation for nonmetricity.
As noted in Section 3, when the generators of any pseudo-orthogonal group $\mathfrak{so}\left(p,q\right)$ are in doubly covariant form, they are always antisymmetric, $\left[G_{\Delta}\right]_{AB}=-\left[G_{\Delta}\right]_{BA}$. The signature of the pseudo-orthogonal metric then enters when we return one index to the raised position, $\left[G_{\Delta}\right]_{\;\;\;B}^{A}=\eta^{AC}\left[G_{\Delta}\right]_{CB}$. We may therefore identify a basis for $\mathfrak{so}\left(3,1\right)$ by finding the antisymmetric connection forms, $\left[\hat{\Gamma}\right]_{\left[AB\right]}$. Starting with the real transformations $\left[\hat{\Gamma}\right]_{\;\;\;B}^{A}$, we may lower $A$ with any convenient nondegenerate real matrix, e.g., $h\simeq\gamma^{0}$ or $\hat{h}=i\hat{\gamma}^{0}\hat{\gamma}^{2}$, since we then take the antisymmetric part and raise with $\eta^{-1}$. Different choices merely assign different names to the same set of Lorentz generators. The simplest choice is the diagonal form $h$.
Writing the covariant matrices $\left[h\hat{\Gamma}\right]_{\left[AB\right]}$ explicitly in the Dirac representation, we find the antisymmetric subset $\left[h\hat{\Gamma}\right]_{\left[AB\right]}\in\left\{ih\gamma^{2},h\sigma^{01},h\sigma^{03},h\sigma^{31},ih\gamma_{5}\gamma^{2},h\gamma_{5}\right\}$. Raising indices with $\eta^{AB}\equiv diag\left(-1,1,1,1\right)$ gives a real basis $\hat{\Gamma}^{\Delta_{a}}$ for $\mathfrak{so}\left(3,1\right)$:
$i\eta h\gamma^{2}=\left(\begin{array}[]{cc}&-\sigma^{1}\\ i\sigma^{2}\end{array}\right),\eta h\hat{\gamma}_{5}\hat{\gamma}^{2}=\left(\begin{array}[]{cc}\sigma^{1}\\ &-i\sigma^{2}\end{array}\right),\eta h\hat{\gamma}_{5}=\left(\begin{array}[]{cc}&-\sigma^{3}\\ -1\end{array}\right)$
$\eta h\sigma^{01}=\left(\begin{array}[]{cc}&-i\sigma^{2}\\ -\sigma^{1}\end{array}\right),\eta h\sigma^{03}=\left(\begin{array}[]{cc}&-1\\ -\sigma^{3}\end{array}\right),\eta h\sigma^{31}=\left(\begin{array}[]{cc}\sigma^{1}\\ &i\sigma^{2}\end{array}\right)$
It is straightforward to check that these span $\mathfrak{so}\left(3,1\right)$. In terms of the usual boost and rotation generators, $\left[K^{i}\right]_{\;\;\;b}^{a}=\delta_{0}^{a}\delta_{b}^{i}+\delta^{ia}\delta_{0b}$, $\left[J^{i}\right]_{\;\;\;b}^{a}=\varepsilon_{\quad b}^{ia}$ respectively
$\begin{array}[]{ccccccc}K_{x}&=&\frac{1}{2}\left(\eta h\hat{\gamma}_{5}\hat{\gamma}^{2}+\eta h\sigma^{31}\right)&&J_{x}&=&\frac{1}{2}\left(\eta h\sigma^{01}-i\eta h\gamma^{2}\right)\\ K_{y}&=&-\frac{1}{2}\left(\eta h\hat{\gamma}_{5}+\eta h\sigma^{03}\right)&&J_{y}&=&\frac{1}{2}\left(\eta h\sigma^{03}-\eta h\hat{\gamma}_{5}\right)\\ K_{z}&=&-\frac{1}{2}\left(i\eta h\gamma^{2}+\eta h\sigma^{01}\right)&&J_{z}&=&\frac{1}{2}\left(\eta h\sigma^{31}-\eta h\hat{\gamma}_{5}\hat{\gamma}^{2}\right)\end{array}$
The remaining 10 combinations
$\eta^{AC}\left[h\hat{\Gamma}\right]_{\left(CB\right)}$ $\in$ $\hat{\Gamma}^{\Delta_{s}}\left\{h1,h\gamma^{0},ih\sigma^{12},ih\sigma^{23},h\gamma_{5}\gamma^{1},h\gamma_{5}\gamma^{3},h\gamma^{1}h\gamma^{3},ih\sigma^{02},h\gamma_{5}\gamma^{0}\right\}$
are symmetric, giving generators
$\eta h1=\left(\begin{array}[]{cc}-\sigma^{3}\\ &-1\end{array}\right),\eta h\gamma^{0}=\left(\begin{array}[]{cc}-\sigma^{3}\\ &1\end{array}\right),\eta h\gamma^{1}=\left(\begin{array}[]{cc}&-i\sigma^{2}\\ -\sigma^{1}\end{array}\right),\eta h\gamma^{3}=\left(\begin{array}[]{cc}&-1\\ \sigma^{3}\end{array}\right),$
$i\eta h\sigma^{02}=\left(\begin{array}[]{cc}&-\sigma^{1}\\ -i\sigma^{2}\end{array}\right),i\eta h\sigma^{12}=\left(\begin{array}[]{cc}-1\\ &-\sigma^{3}\end{array}\right),i\eta h\sigma^{23}=\left(\begin{array}[]{cc}-i\sigma^{2}\\ &-\sigma^{1}\end{array}\right),$
$\eta h\hat{\gamma}_{5}\hat{\gamma}^{1}=\left(\begin{array}[]{cc}i\sigma^{2}\\ &-\sigma^{1}\end{array}\right),\eta h\hat{\gamma}_{5}\hat{\gamma}^{3}=\left(\begin{array}[]{cc}1\\ &-\sigma^{3}\end{array}\right),\eta h\hat{\gamma}_{5}\hat{\gamma}^{2}=\left(\begin{array}[]{cc}\sigma^{1}\\ &-i\sigma^{2}\end{array}\right)$
Varying a real linear combination of these will give our source for nonmetricity.
The full connection is therefore $\boldsymbol{\Sigma}=\mathbf{b}_{\Delta}\Gamma^{\Delta}=\mathbf{b}_{\Delta_{s}}\Gamma^{\Delta_{s}}+\mathbf{b}_{\Delta_{a}}\Gamma^{\Delta_{a}}$. Explicitly
$\Omega_{\;\;\;Bc}^{A}=b_{c\Delta_{a}}\left[\Gamma^{\Delta_{a}}\right]_{\;\;\;B}^{A}$ $=$ $\left[ib_{c2}\eta h\gamma^{2}+b_{c01}\eta h\sigma^{01}+b_{c03}\eta h\sigma^{03}+b_{c31}\eta h\sigma^{31}+ib_{c52}\eta h\gamma_{5}\gamma^{2}+b_{c5}\eta h\gamma_{5}\right]_{\;\;\;B}^{A}$
$Q_{\;\;\;Bc}^{A}=b_{c\Delta_{s}}\left[\Gamma^{\Delta_{s}}\right]_{\;\;\;B}^{A}$ $=$ $\left[b_{c}\eta h1\right.+b_{c0}\eta h\gamma^{0}+b_{c1}\eta h\gamma^{1}+b_{c3}\eta h\gamma^{3}+ib_{c02}\eta h\sigma^{02}+ib_{c12}\eta h\sigma^{12}+ib_{c23}\eta h\sigma^{23}$ (29)
$+b_{50}\eta h\gamma_{5}\gamma^{0}+b_{c51}\eta h\gamma_{5}\gamma^{1}+\left.b_{c53}\eta h\gamma_{5}\gamma^{3}\right]_{\;\;\;B}^{A}$
Varying $\left(b_{cij},b_{c2},b_{c52},b_{c5}\right)_{i=0,1,3}$ will give the source for torsion while varying $\left(b_{c},b_{ci},b_{c2i},b_{c5i}\right)_{i=0,1,3}$ will give the source for nonmetricity.
5 The field equations
Writing the gravitational variation (13) in the $\mathfrak{Cl}\left(2,2\right)$ expansion, we use the antisymmetry of $\left[\Gamma^{\Delta_{a}}\right]^{AB}$ and the symmetry of $\left[\Gamma^{\Delta_{s}}\right]^{AB}$ to write the products as traces.
$\delta\Omega_{\qquad c}^{ab}\mathscr{T}_{\;\;\;ab}^{c}$ $=$ $b_{c\Delta_{a}}\left[\Gamma^{\Delta_{a}}\right]^{AB}\mathscr{T}_{\;\;\;AB}^{c}=-b_{c\Delta_{a}}tr\left(\Gamma^{\Delta_{a}}\mathscr{T}^{c}\right)$
$-\frac{1}{4}\delta Q_{\quad a}^{bc}\mathscr{Q}_{\;\;\;bc}^{a}$ $=$ $-\frac{1}{4}b_{c\Delta_{s}}\left[\Gamma^{\Delta_{s}}\right]^{AB}\mathscr{Q}_{\;\;\;AB}^{c}=-\frac{1}{4}b_{c\Delta_{s}}tr\left(\Gamma^{\Delta_{s}}\mathscr{Q}^{c}\right)$
where $\mathscr{T}_{\;\;\;ab}^{c}$ and $\mathscr{Q}_{\;\;\;bc}^{a}$ are given by Eqs.(9) and (11), respectively. The field equations follow from the variation $\delta S_{Grav}+\delta S_{D}=0$.
$-\kappa\delta b_{c\Delta_{a}}tr\left(\Gamma^{\Delta_{a}}\mathscr{T}^{c}\right)-\frac{\kappa}{4}\delta b_{c\Delta_{s}}tr\left(\Gamma^{\Delta_{s}}\mathscr{Q}_{c}\right)$ $=$ $\alpha Re\left(i\psi^{\dagger}h\gamma^{c}\delta b_{\Delta c}\hat{\Gamma}^{\Delta}\psi\right)$ (30)
Next, we expand each side explicitly.
5.1 Gravitational interaction
Substituting the explicit connection from Eqs.(29) into Eq.(31), We find the gravitational contribution to the field equation by expanding the sums on $\Delta_{a}$ and $\Delta_{s}$. Notice that once we take the traces, the expressions are real components of Lorentzian matrices, so we revert to lower case Latin indices. Writing out the full sum,
$\delta\mathcal{L}_{Grav}$ $=$ $\kappa\delta\Omega_{\quad c}^{AB}\mathcal{T}_{\;\;\;AB}^{c}-\frac{\kappa}{4}\delta Q^{ABc}\mathscr{Q}_{c\left(AB\right)}$ (31)
$=$ $-\kappa\delta b_{c2}tr\left[i\eta h\gamma^{2}\mathcal{\mathscr{T}}^{c}\right]-\kappa\delta b_{c01}tr\left[\eta h\sigma^{01}\mathcal{\mathscr{T}}^{c}\right]-\kappa\delta b_{c03}tr\left[\eta h\sigma^{03}\mathcal{\mathscr{T}}^{c}\right]$
$-\kappa\delta b_{c13}tr\left[\eta h\sigma^{13}\mathcal{\mathscr{T}}^{c}\right]-\kappa\delta b_{c52}tr\left[i\eta h\gamma_{5}\gamma^{2}\mathcal{\mathscr{T}}^{c}\right]-\kappa\delta b_{c5}tr\left[\eta h\gamma_{5}\mathcal{\mathscr{T}}^{c}\right]$
$-\frac{\kappa}{4}\delta b_{c}tr\left[\eta h1\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c0}tr\left[\eta h\gamma^{0}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c1}tr\left[\eta h\gamma^{1}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c3}tr\left[\eta h\gamma^{3}\mathcal{\mathscr{Q}}^{c}\right]$
$-\frac{\kappa}{4}\delta b_{c02}tr\left[i\eta h\sigma^{02}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c12}tr\left[i\eta h\sigma^{12}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c23}tr\left[i\eta h\sigma^{23}\mathcal{\mathscr{Q}}^{c}\right]$
$-\frac{\kappa}{4}\delta b_{c50}tr\left[\eta h\gamma_{5}\gamma^{0}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c51}tr\left[\eta h\gamma_{5}\gamma^{1}\mathcal{\mathscr{Q}}^{c}\right]-\frac{\kappa}{4}\delta b_{c53}tr\left[\eta h\gamma_{5}\gamma^{3}\mathcal{\mathscr{Q}}^{c}\right]$
we then compute each of the 16 traces. The result is a collection of linear combinations of components of $\mathcal{T}_{\;\;\;ab}^{c}$. We find the explicit combinations by carrying out the traces in the Dirac representation.
For the torsion, $\mathcal{T}_{\quad ab}^{c}$ these we find
$-\left[i\eta h\gamma^{2}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $\mathcal{T}_{\quad 30}^{c}+\mathcal{T}_{\quad 21}^{c}-\mathcal{T}_{\quad 12}^{c}-\mathcal{T}_{\quad 03}^{c}$
$-\left[\eta h\sigma^{01}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $\mathcal{T}_{\quad 30}^{c}-\mathcal{T}_{\quad 21}^{c}+\mathcal{T}_{\quad 12}^{c}-\mathcal{T}_{\quad 03}^{c}$
$-\left[i\eta h\gamma_{5}\gamma^{2}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\mathcal{T}_{\quad 10}^{c}+\mathcal{T}_{\quad 01}^{c}+\mathcal{T}_{\quad 32}^{c}-\mathcal{T}_{\quad 23}^{c}$
$-\left[\eta h\sigma^{31}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\mathcal{T}_{\quad 10}^{c}+\mathcal{T}_{\quad 01}^{c}-\mathcal{T}_{\quad 32}^{c}+\mathcal{T}_{\quad 23}^{c}$
$-\left[\eta h\gamma_{5}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $\mathcal{T}_{\quad 20}^{c}-\mathcal{T}_{\quad 31}^{c}-\mathcal{T}_{\quad 02}^{c}+\mathcal{T}_{\quad 13}^{c}$
$-\left[\eta h\sigma^{03}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $\mathcal{T}_{\quad 20}^{c}+\mathcal{T}_{\quad 31}^{c}-\mathcal{T}_{\quad 02}^{c}-\mathcal{T}_{\quad 13}^{c}$ (32)
The traces for nonmetricity are
$tr\left[\eta h1\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathscr{Q}}_{\quad 00}^{c}+\mathcal{\mathscr{Q}}_{\quad 11}^{c}-\mathcal{\mathscr{Q}}_{\quad 22}^{c}-\mathcal{\mathscr{Q}}_{\quad 33}^{c}$
$tr\left[\eta h\gamma^{0}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathscr{Q}}_{\quad 00}^{c}+\mathcal{\mathscr{Q}}_{\quad 11}^{c}+\mathcal{\mathscr{Q}}_{\quad 22}^{c}+\mathcal{\mathscr{Q}}_{\quad 33}^{c}$
$tr\left[\eta h\gamma^{1}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $-\mathcal{\mathscr{Q}}_{\quad 30}^{c}+\mathcal{\mathscr{Q}}_{\quad 21}^{c}+\mathcal{\mathscr{Q}}_{\quad 12}^{c}-\mathcal{\mathscr{Q}}_{\quad 03}^{c}$
$tr\left[\eta h\gamma^{3}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $-\mathcal{\mathcal{\mathscr{Q}}}_{\quad 20}^{c}-\mathcal{\mathscr{Q}}_{\quad 31}^{c}-\mathcal{\mathscr{Q}}_{\quad 02}^{c}-\mathcal{\mathscr{Q}}_{\quad 13}^{c}$
$tr\left[i\eta h\sigma^{02}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $-\mathcal{\mathscr{Q}}_{\quad 30}^{c}-\mathcal{\mathscr{Q}}_{\quad 12}^{c}-\mathcal{\mathscr{Q}}_{\quad 03}^{c}-\mathcal{\mathscr{Q}}_{\quad 21}^{c}$
$tr\left[i\eta h\sigma^{12}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathscr{Q}}_{\quad 00}^{c}-\mathcal{\mathscr{Q}}_{\quad 11}^{c}-\mathcal{\mathscr{Q}}_{\quad 22}^{c}+\mathcal{\mathscr{Q}}_{\quad 33}^{c}$
$tr\left[i\eta h\sigma^{32}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathscr{Q}}_{\quad 10}^{c}+\mathcal{\mathscr{Q}}_{\quad 01}^{c}+\mathcal{\mathscr{Q}}_{\quad 32}^{c}+\mathcal{\mathscr{Q}}_{\quad 23}^{c}$
$tr\left[\eta h\gamma_{5}\gamma^{0}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathcal{\mathscr{Q}}}_{\quad 20}^{c}-\mathcal{\mathscr{Q}}_{\quad 31}^{c}+\mathcal{\mathscr{Q}}_{\quad 02}^{c}-\mathcal{\mathscr{Q}}_{\quad 13}^{c}$
$tr\left[\eta h\gamma_{5}\gamma^{1}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $\mathcal{\mathscr{Q}}_{\quad 10}^{c}+\mathcal{\mathscr{Q}}_{\quad 01}^{c}-\mathcal{\mathscr{Q}}_{\quad 32}^{c}-\mathcal{\mathscr{Q}}_{\quad 23}^{c}$
$tr\left[\eta h\gamma_{5}\gamma^{3}\mathcal{\mathscr{Q}}^{c}\right]$ $=$ $-\mathcal{\mathscr{Q}}_{\quad 00}^{c}+\mathcal{\mathscr{Q}}_{\quad 11}^{c}-\mathcal{\mathscr{Q}}_{\quad 22}^{c}+\mathcal{\mathscr{Q}}_{\quad 33}^{c}$ (33)
For each $c=0,1,2,3$ these are equated to scalars built from Dirac spinors.
5.2 Spinor interaction
The sum over $\Delta$ in the spinor contribution to Eq.(30) gives
$\delta\mathcal{L}_{interaction}$ $=$ $-\alpha Re\left(\psi^{\dagger}\left(i\delta b_{a01}h\gamma^{a}\eta h\sigma^{01}+i\delta b_{a03}h\gamma^{a}\eta h\sigma^{03}+i\delta b_{a13}h\gamma^{a}\eta h\sigma^{13}\right)\psi\right)$ (34)
$-\alpha Re\left(\psi^{\dagger}\left(i\delta b_{a2}ih\gamma^{a}\eta h\gamma^{2}+i\delta b_{a52}ih\gamma^{a}\eta h\gamma_{5}\gamma^{2}+i\delta b_{a5}h\gamma^{a}\eta h\gamma_{5}\right)\psi\right)$
$-\alpha\delta b_{c}Re\psi^{\dagger}\left(ih\gamma^{c}\eta h1\right)\psi-\alpha Re\left(\psi^{\dagger}\left(i\delta b_{a0}h\gamma^{a}\eta h\gamma^{0}+i\delta b_{a1}h\gamma^{a}\eta h\gamma^{1}+i\delta b_{a3}h\gamma^{a}\eta h\gamma^{3}\right)\psi\right)$
$-\alpha\left(Re\psi^{\dagger}\left(i\delta b_{a50}h\gamma^{a}\eta h\gamma_{5}\gamma^{0}+i\delta b_{a51}h\gamma^{a}\eta h\gamma_{5}\gamma^{1}+i\delta b_{a53}h\gamma^{a}\eta h\gamma_{5}\gamma^{3}\right)\psi\right)$
$-\alpha Re\left(\psi^{\dagger}\left(i\delta b_{a02}ih\gamma^{a}\eta h\sigma^{02}+i\delta b_{a12}ih\gamma^{a}\eta h\sigma^{12}+i\delta b_{a32}ih\gamma^{a}\eta h\sigma^{32}\right)\psi\right)$
where there remains a sum on $a=0,1,2,3$. We list the torsion coefficients $\left\{b_{a01},b_{a03},b_{a13},b_{a2},b_{a52},b_{a5}\right\}$ first.
The breaking of Lorentz symmetry by the general linear group disrupts the usual systematic index notation so that each term must be computed individually. The computation of each term is straightforward. Throughout we choose the spinor components as
$\psi$ $=$ $\left(\begin{array}[]{c}\mu\\ \nu\\ \rho\\ \sigma\end{array}\right),\psi^{\dagger}=\left(\bar{\mu},\bar{\nu},\bar{\rho},\bar{\sigma}\right)$
For example, the third term in Eq.(34), $\alpha b_{31}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{1}\psi\right)$ becomes
$-\alpha Re\left(\psi^{\dagger}b_{31}ih\gamma^{3}\eta h\gamma^{1}\psi\right)$ $=$ $-\alpha b_{31}Re\left(i\psi^{\dagger}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\left(\begin{array}[]{cc}&\sigma^{3}\\ -\sigma^{3}\end{array}\right)\left(\begin{array}[]{cc}-\sigma^{3}\\ &1\end{array}\right)\left(\begin{array}[]{cc}1\\ &-1\end{array}\right)\left(\begin{array}[]{cc}&\sigma^{1}\\ -\sigma^{1}\end{array}\right)\psi\right)$
$=$ $-i\alpha b_{31}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)$
This simplification is carried out for each of the 64 terms in Eq.(34). The resulting Dirac scalars are given in the Appendix Appendix: Spinor couplings.
5.3 Field equations
This Subsection contains our principal results.
The field equations $\delta\mathcal{L}_{Grav}=-\delta\mathcal{L}_{Interaction}$ follow by equating the variation coefficients $\delta b_{cA}$. For the torsion these become the 24 equations
$\left[\eta hi\gamma^{2}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}i\eta h\gamma^{2}\psi\right)$
$\left[\eta h\sigma^{01}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}\eta h\sigma^{01}\psi\right)$
$\left[\eta h\gamma_{5}i\gamma^{2}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}i\eta h\gamma_{5}\gamma^{2}\psi\right)$
$\left[\eta h\sigma^{31}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}\eta h\sigma^{31}\psi\right)$
$\left[\eta h\gamma_{5}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}\eta h\gamma_{5}\psi\right)$
$\left[\eta h\sigma^{03}\right]_{\;\;\;b}^{a}\mathcal{T}_{\quad a}^{cb}$ $=$ $-\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{c}\eta h\sigma^{03}\psi\right)$
This is the $\mathfrak{spin}\left(2,2\right)$ decomposition of the torsion, with each independent projection sourced by a different current. The $\mathfrak{spin}\left(2,2\right)$ decomposition of the nonmetricity is comprised of the remaining 10 equations.
$\left[\eta h\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\psi\right)$
$\left[\eta h\gamma^{0}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma^{0}\psi\right)$
$\left[\eta h\gamma^{1}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma^{1}\psi\right)$
$\left[\eta h\gamma^{3}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma^{3}\psi\right)$
$\left[\eta h\gamma_{5}\gamma^{0}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma_{5}\gamma^{0}\psi\right)$
$\left[\eta h\gamma_{5}\gamma^{1}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma_{5}\gamma^{1}\psi\right)$
$\left[\eta h\gamma_{5}\gamma^{3}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}\eta h\gamma_{5}\gamma^{3}\psi\right)$
$\left[i\eta h\sigma^{02}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}i\eta h\sigma^{02}\psi\right)$
$\left[i\eta h\sigma^{12}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}i\eta h\sigma^{12}\psi\right)$
$\left[i\eta h\sigma^{23}\right]_{\;\;\;b}^{c}\mathscr{Q}_{\quad c}^{ab}$ $=$ $-\frac{4\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{a}i\eta h\sigma^{23}\psi\right)$
The sources on the right are built from the components of $\psi$, but because of the factor of $\eta$ they do not correspond directly to the usual Dirac currents. To see the detailed form of the sources, we replace each component on the right with their scalar expansions. These are listed in the Appendix Appendix: Spinor couplings. We replace the corresponding linear combinations on the left with the expansions in Eqs.(32) and (33), and it is then easy to solve for the individual components $\mathcal{T}_{\;\;\;bc}^{a}$ of the torsion and $\mathscr{Q}_{\;\;\;bc}^{a}$ of the nonmetricity.
Write the real, imaginary, and diagonal parts of complex products as the real numbers:
$R_{\bar{\alpha}\beta}$ $=$ $\frac{1}{2}\left(\bar{\alpha}\beta+\bar{\beta}\alpha\right)$
$I_{\bar{\alpha}\beta}$ $=$ $\frac{1}{2i}\left(\bar{\alpha}\beta-\bar{\beta}\alpha\right)$
$D_{\bar{\alpha}\alpha}$ $=$ $\bar{\alpha}\alpha$
Then the results for torsion are:
$\mathcal{T}_{\;\;\;ab}^{0}$ $=$ $\frac{\alpha}{\kappa}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&-I_{\bar{\nu}\rho}&-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\nu}\rho}&0&-I_{\bar{\rho}\sigma}\\ 0&I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{1}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\sigma}+I_{\bar{\mu}\rho}&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0\\ -I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}\\ -I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&0&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{2}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&R_{\bar{\mu}\rho}-R_{\bar{\nu}\sigma}&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&0&\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}\\ R_{\bar{\mu}\nu}+R_{\bar{\rho}\sigma}&-\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&0&-\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)\\ -\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)&R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{3}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&0&-\left(I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\right)\\ I_{\bar{\mu}\sigma}-I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&-\left(I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}\right)\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}&0\end{array}\right)$
The results for nonmetricity are:
$\mathscr{Q}_{\quad ab}^{0}$ $=$ $\frac{4\alpha}{\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\mu}\nu}&I_{\bar{\mu}\rho}&I_{\bar{\mu}\sigma}\\ I_{\bar{\mu}\nu}&0&0&0\\ I_{\bar{\mu}\rho}&0&0&0\\ I_{\bar{\mu}\sigma}&0&0&0\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{1}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}2I_{\bar{\mu}\sigma}&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&-2I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}\\ I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0&2I_{\bar{\nu}\rho}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}&2I_{\bar{\mu}\sigma}\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{2}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2R_{\bar{\mu}\sigma}&R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&-2R_{\bar{\nu}\rho}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}\\ R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&2R_{\bar{\nu}\rho}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}\\ 2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&-2R_{\bar{\mu}\sigma}\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{3}$ $=$ $-\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&0&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\\ I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&-2I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&2I_{\bar{\nu}\sigma}\end{array}\right)$
Note that the nonmetricity has vanishing trace, $\eta^{ab}\mathscr{Q}_{\quad ab}^{c}=0$, showing that the Dirac equation does not couple to the Weyl vector, in agreement with [29].
6 Special cases
We gain some insight by looking at restricted spinors. For a spin-up electron at rest, $\psi$ reduces to $\psi_{e0}=\left(\mu,0,0,0\right)$. Then all components of the torsion vanish $\mathcal{T}_{\;\;\;ab}^{0}=\mathcal{T}_{\;\;\;ab}^{1}=\mathcal{T}_{\;\;\;ab}^{3}=0$ except for
$\mathcal{T}_{\;\;\;ab}^{2}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&&&D_{\bar{\mu}\mu}\\ &0\\ &&0\\ -D_{\bar{\mu}\mu}&&&0\end{array}\right)$
Similarly, a spin up positron at rest with $\psi_{p0}=\left(0,0,0,\sigma\right)$ produces torsion given by $\mathcal{T}_{\;\;\;ab}^{0}=\mathcal{T}_{\;\;\;ab}^{1}=\mathcal{T}_{\;\;\;ab}^{3}=0$ and
$\mathcal{T}_{\;\;\;ab}^{2}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&&&-D_{\bar{\sigma}\sigma}\\ &0\\ &&0\\ D_{\bar{\sigma}\sigma}&&&0\end{array}\right)$
These differ only by the replacement $\mu\rightarrow\sigma$ and an overall sign. The simple relationship between particle and antiparticle will be maintained by the $SO\left(3,1\right)$ subgroup since interchange of past and future light cones is a Lorentz transformation.
Since nonmetricity breaks Lorentz invariance, the particle/antiparticle cases may differ. However, the nonmetricity arising from a spin-up electron is $\mathscr{Q}_{\quad ab}^{0}=\mathscr{Q}_{\quad ab}^{1}=\mathscr{Q}_{\quad ab}^{3}=0$ and
$\mathscr{Q}_{\quad ab}^{2}$ $=$ $\frac{4\alpha}{\kappa}\left(\begin{array}[]{cccc}0&&&D_{\bar{\mu}\mu}\\ &0\\ &&0\\ D_{\bar{\mu}\mu}&&&0\end{array}\right)$
while that of a spin-up positron at rest is similar, $\mathscr{Q}_{\quad ab}^{0}=\mathscr{Q}_{\quad ab}^{1}=\mathscr{Q}_{\quad ab}^{3}=0$, and
$\mathscr{Q}_{\quad ab}^{2}$ $=$ $\frac{4\alpha}{\kappa}\left(\begin{array}[]{cccc}0&&&D_{\bar{\sigma}\sigma}\\ &0\\ &&0\\ D_{\bar{\sigma}\sigma}&&&0\end{array}\right)$
Other special cases follow easily.
Acknowledgements:
The author thanks Matthew Pelligrini and Joshua Knobloch for useful discussions.
Appendix: Spinor couplings
The interaction terms may be written as $S_{interaction}=-\alpha\int Re\left(i\psi^{\dagger}\left(h\gamma^{a}b_{Aa}\eta h\hat{\Gamma}^{A}\right)\psi\right)$. When we expand the connection this gives 64 scalars, divided by symmetry, 24 for torsion sources and 40 for nonmetric sources. Here we list the resulting scalars.
Spinor scalars for torsion
The spinor scalars required for torsion sources are
$\begin{array}[]{ccccccccc}\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}i\eta h\gamma^{2}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}i\eta h\gamma^{2}\psi\right)&=&\frac{\alpha}{\kappa}\left(-\bar{\mu}\mu-\bar{\nu}\nu-\bar{\rho}\rho+\bar{\sigma}\sigma\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\sigma^{01}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\sigma^{01}\psi\right)&=&\frac{\alpha}{\kappa}\left(-\bar{\mu}\mu+\bar{\nu}\nu+\bar{\rho}\rho+\bar{\sigma}\sigma\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}i\eta h\gamma_{5}\gamma^{2}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}i\eta h\gamma_{5}\gamma^{2}\psi\right)&=&\frac{\alpha}{\kappa}\left(\bar{\mu}\rho+\bar{\rho}\mu\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\sigma^{31}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\sigma^{31}\psi\right)&=&-\frac{\alpha}{\kappa}\left(\bar{\nu}\sigma+\bar{\sigma}\nu\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma_{5}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma_{5}\psi\right)&=&\frac{\alpha}{\kappa}\left(\bar{\rho}\sigma+\bar{\sigma}\rho\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\sigma^{03}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\sigma^{03}\psi\right)&=&\frac{\alpha}{\kappa}\left(\bar{\mu}\nu+\bar{\nu}\mu\right)\\ \\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}i\eta h\gamma^{2}\psi\right)&=&0&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}i\eta h\gamma^{2}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\sigma^{01}\psi\right)&=&0&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\sigma^{01}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}i\eta h\gamma_{5}\gamma^{2}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}i\eta h\gamma_{5}\gamma^{2}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\sigma^{31}\psi\right)&=&-\frac{i\alpha}{\kappa}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\sigma^{31}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\mu}\sigma-\bar{\sigma}\mu\right)\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma_{5}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma_{5}\psi\right)&=&0\\ \frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\sigma^{03}\psi\right)&=&\frac{i\alpha}{\kappa}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)&&&&\frac{\alpha}{\kappa}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\sigma^{03}\psi\right)&=&0\end{array}$
Spinor scalars for nonmetricity
The spinor scalars required for nonmetricity sources are
$\begin{array}[]{ccccccc}\alpha\delta b_{0}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\psi\right)&=&0&&\alpha\delta b_{2}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\psi\right)&=&\alpha\delta b_{2}\left(\bar{\nu}\rho+\bar{\rho}\nu\right)\\ \alpha\delta b_{00}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma^{0}\psi\right)&=&0&&\alpha\delta b_{20}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma^{0}\psi\right)&=&\alpha\delta b_{20}\left(\bar{\mu}\sigma+\bar{\sigma}\mu\right)\\ \alpha\delta b_{01}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma^{1}\psi\right)&=&i\alpha\delta b_{01}\left(\bar{\sigma}\mu-\bar{\mu}\sigma\right)&&\alpha\delta b_{21}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma^{1}\psi\right)&=&\alpha\delta b_{21}\left(\bar{\mu}\mu-\bar{\nu}\nu+\bar{\rho}\rho+\bar{\sigma}\sigma\right)\\ \alpha\delta b_{03}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma^{3}\psi\right)&=&-i\alpha\delta b_{03}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)&&\alpha\delta b_{23}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma^{3}\psi\right)&=&-\alpha\delta b_{23}Re\left(\bar{\mu}\nu+\bar{\nu}\mu\right)\\ \alpha\delta b_{050}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma_{5}\gamma^{0}\psi\right)&=&i\alpha\delta b_{050}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)&&\alpha\delta b_{250}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma_{5}\gamma^{0}\psi\right)&=&-\alpha\delta b_{250}\left(\bar{\rho}\sigma+\bar{\sigma}\rho\right)\\ \alpha\delta b_{051}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma_{5}\gamma^{1}\psi\right)&=&i\alpha\delta b_{051}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)&&\alpha\delta b_{251}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma_{5}\gamma^{1}\psi\right)&=&-\alpha\delta b_{251}\left(\bar{\mu}\rho+\bar{\rho}\mu\right)\\ \alpha\delta b_{053}Re\left(\psi^{\dagger}ih\gamma^{0}\eta h\gamma_{5}\gamma^{3}\psi\right)&=&0&&\alpha\delta b_{253}Re\left(\psi^{\dagger}ih\gamma^{2}\eta h\gamma_{5}\gamma^{3}\psi\right)&=&\alpha\delta b_{253}\left(\bar{\nu}\rho+\bar{\rho}\nu\right)\\ \alpha\delta b_{002}Re\left(\psi^{\dagger}ih\gamma^{0}i\eta h\sigma^{02}\psi\right)&=&-i\alpha\delta b_{002}\left(\bar{\mu}\sigma-\bar{\sigma}\mu\right)&&\alpha\delta b_{202}Re\left(\psi^{\dagger}ih\gamma^{2}i\eta h\sigma^{02}\psi\right)&=&\alpha\delta b_{202}\left(\bar{\mu}\mu+\bar{\nu}\nu-\bar{\rho}\rho+\bar{\sigma}\sigma\right)\\ \alpha\delta b_{012}Re\left(\psi^{\dagger}ih\gamma^{0}i\eta h\sigma^{12}\psi\right)&=&0&&\alpha\delta b_{212}Re\left(\psi^{\dagger}ih\gamma^{2}i\eta h\sigma^{12}\psi\right)&=&\alpha\delta b_{212}\left(\bar{\mu}\sigma+\bar{\sigma}\mu\right)\\ \alpha\delta b_{023}Re\left(\psi^{\dagger}ih\gamma^{0}i\eta h\sigma^{23}\psi\right)&=&-i\alpha\delta b_{023}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)&&\alpha\delta b_{223}Re\left(\psi^{\dagger}ih\gamma^{2}i\eta h\sigma^{23}\psi\right)&=&\alpha\delta b_{223}\left(\bar{\nu}\sigma+\bar{\sigma}\nu\right)\\ \\ \alpha\delta b_{1}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\psi\right)&=&-i\alpha\delta b_{1}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)&&\alpha\delta b_{3}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\psi\right)&=&i\alpha\delta b_{3}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)\\ \alpha\delta b_{10}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma^{0}\psi\right)&=&i\alpha\delta b_{10}\left(\bar{\mu}\sigma-\bar{\sigma}\mu\right)&&\alpha\delta b_{30}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{0}\psi\right)&=&i\alpha\delta b_{30}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)\\ \alpha\delta b_{11}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma^{1}\psi\right)&=&0&&\alpha\delta b_{31}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{1}\psi\right)&=&i\alpha\delta b_{31}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)\\ \alpha\delta b_{13}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma^{3}\psi\right)&=&-i\alpha\delta b_{13}\left(\bar{\mu}\nu-\bar{\nu}\mu\right)&&\alpha\delta b_{33}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{3}\psi\right)&=&0\\ \alpha\delta b_{150}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma_{5}\gamma^{0}\psi\right)&=&-i\alpha\delta b_{150}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)&&\alpha\delta b_{350}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma_{5}\gamma^{0}\psi\right)&=&0\\ \alpha\delta b_{151}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma_{5}\gamma^{1}\psi\right)&=&-i\alpha\delta b_{151}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)&&\alpha\delta b_{351}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma_{5}\gamma^{1}\psi\right)&=&-i\alpha\delta b_{351}\left(\bar{\mu}\sigma-\bar{\sigma}\mu\right)\\ \alpha\delta b_{153}Re\left(\psi^{\dagger}ih\gamma^{1}\eta h\gamma_{5}\gamma^{3}\psi\right)&=&-i\alpha\delta b_{153}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)&&\alpha\delta b_{353}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma_{5}\gamma^{3}\psi\right)&=&-i\alpha\delta b_{353}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)\\ \alpha\delta b_{102}Re\left(\psi^{\dagger}ih\gamma^{1}i\eta h\sigma^{02}\psi\right)&=&0&&\alpha\delta b_{302}Re\left(\psi^{\dagger}ih\gamma^{3}i\eta h\sigma^{02}\psi\right)&=&i\alpha\delta b_{302}\left(\bar{\rho}\sigma-\bar{\sigma}\rho\right)\\ \alpha\delta b_{112}Re\left(\psi^{\dagger}ih\gamma^{1}i\eta h\sigma^{12}\psi\right)&=&i\alpha\delta b_{112}\left(\bar{\mu}\sigma-\bar{\sigma}\mu\right)&&\alpha\delta b_{312}Re\left(\psi^{\dagger}ih\gamma^{3}i\eta h\sigma^{12}\psi\right)&=&-i\alpha\delta b_{312}\left(\bar{\nu}\sigma-\bar{\sigma}\nu\right)\\ \alpha\delta b_{123}Re\left(\psi^{\dagger}ih\gamma^{1}i\eta h\sigma^{23}\psi\right)&=&-i\alpha\delta b_{123}\left(\bar{\mu}\rho-\bar{\rho}\mu\right)&&\alpha\delta b_{323}Re\left(\psi^{\dagger}ih\gamma^{3}i\eta h\sigma^{23}\psi\right)&=&i\alpha\delta b_{323}\left(\bar{\nu}\rho-\bar{\rho}\nu\right)\end{array}$